OpenStudy (anonymous):

Gold stays solid at Fahrenheit temperatures below 1945.4°. Determine (in terms of an inequality) those Celsius temperatures for which gold stays solid. Using the formula f=9/5c+32

OpenStudy (anonymous):

F=9/5C+32 1945.4=9/5C+32 1945.4-32=9/5c+32-32 1913.4=9/5c 1913.4÷9/5 or 1.8=c 1063=c This is what I came up with but it isn't in the terms of inequality :-s